maddie123:

If d = 8, what is the value of the expression 4 + d/2? A. 4 B. 6 C. 8 D. 12 E. 14

NerdyLizard:

Ok, have you tried inputting the d value given into the equation?

maddie123:

yes

NerdyLizard:

It would make 4+(8)/2. Good!

maddie123:

i just dont get it please help

NerdyLizard:

Your next step would be to solve it. So, 8/2=4 and 4+4=

maddie123:

ok

NerdyLizard:

You must remember pemdas.

maddie123:

ok

NerdyLizard:

Paranthesis, Exponent, Multiply, Divide, Add, Subtract.

NerdyLizard:

With that mnemonic you can remember the order.

maddie123:

is it 8?

bizaardvark:

If d=8 then you would want to plug in 8 into the expression: 4 + (8)2 4 + 4 = ?

NerdyLizard:

Righto Maddie! Good job!

maddie123:

oh thanks